OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

sorry no :/

OpenStudy (skullpatrol):

Have you seen the slope intercept form: $$\Huge y=mx +b$$

OpenStudy (anonymous):

yes that i have seen

OpenStudy (conqueror):

Those are Skullpatrol's favorite two words on every post.

OpenStudy (skullpatrol):

Ok, so change your equation: 3x – 2y = –6 into that form.

OpenStudy (anonymous):

so would the final equation be Y=-3/2X+8

OpenStudy (skullpatrol):

Great work :-)

OpenStudy (anonymous):

thank you for the help sir :3

OpenStudy (skullpatrol):

Now your question asks for "What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?"

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

You have the slope: -3/2 and a given point (4,2) can you give the equation of that line?

OpenStudy (anonymous):

Y=-3/2X+8?

OpenStudy (skullpatrol):

That is the equation of the original line. You want a new equation: slope: -3/2 and a given point (4,2)

OpenStudy (skullpatrol):

You need to calculate the value of b for the new line using y = (-3/2)x + b using (4,2) for (x, y)

OpenStudy (anonymous):

my four choices are Y=-3/2x+8 Y=3/2x-4 y=1/2x and y=2/3x-4

OpenStudy (loser66):

another way to calculate